# Exploring the range of impacts of helium in the spectra of double detonation models for Type Ia supernovae

### v1

@@ Abstract
In the double detonation scenario the ignition of a surface He detonation on a sub-Chandrasekhar mass white dwarf leads to a secondary core detonation. Double detonation models have shown promise for explaining Type Ia supernovae (SNe Ia) with a variety of luminosities. A key feature of such models is unburnt He in the ejecta, which can show significant variation in both its mass and velocity distribution. Many previous radiative transfer simulations for double detonation models have neglected treatment of non-thermal ionization and excitation, preventing them from robustly assessing whether He spectral features are expected to form. We present a non local thermodynamic equilibrium radiative transfer simulation, including treatment for non-thermal electrons, for a double detonation model with a modest mass of He (0.04 ) ejected at reasonably low velocities (). Despite our simulation predicting no clear optical He features, a strong and persistent He i 10830 Å absorption feature forms that is significantly blended with the spectral contribution of Mg ii 10927 Å. For some normal SNe Ia the Mg feature shows an extended blue wing, previously attributed to C i, however the simulated He feature shows its strongest absorption at wavelengths consistent with this wing. We therefore suggest this extended wing is instead a spectral signature of He. The He feature predicted by this particular model is too strong and persistent to be consistent with normal SNe Ia, however, this motivates further work to use this observable signature to test the parameter space for double detonation models.
@@
Type Ia supernovae (SNe Ia) are the thermonuclear explosions of at least one white dwarf (WD) in a close binary system. However, the nature of the progenitor systems and explosion mechanisms that produce SNe Ia are still uncertain. The double detonation is a widely discussed sub-Chandrasekhar (sub-) explosion mechanism which has the potential to explain SNe Ia with a variety of luminosities. This scenario involves the ignition of a He detonation in a surface He layer (shell) on a carbon-oxygen (CO) WD (see, e.g., the study of He ignition by Glasner et al., 2018). The shock from the initial He detonation then results in the ignition of a secondary core detonation that entirely disrupts the WD. The massive He shells ( ) adopted for early double detonation models (e.g. Nomoto 1980, 1982; Taam 1980; Livne 1990; Woosley & Weaver 1994; Hoeflich & Khokhlov 1996; Nugent et al. 1997) led to an over abundance of iron group elements in the outer ejecta produced by the He shell detonation, resulting in light curves and spectra incompatible with observations of normal SNe Ia. However, more recent studies suggest models with significantly lower He shell masses (< 0.1 ) can result in a secondary core detonation (Bildsten et al., 2007; Shen & Bildsten, 2009; Fink et al., 2010; Shen et al., 2010; Shen & Moore, 2014; Tanikawa et al., 2018, 2019; Gronow et al., 2020; Gronow et al., 2021; Boos et al., 2021; Pakmor et al., 2022; Shen et al., 2024), and could lead to better agreement with observed SNe Ia (Kromer et al., 2010; Woosley & Kasen, 2011). Simulations of such lower He shell mass double detonation models can also predict light curves and spectra broadly consistent with normal SNe Ia (Townsley et al. 2019; Polin et al. 2019; Shen et al. 2021a; Pollin et al. 2024). However, Collins et al. (2022) demonstrate this is not necessarily the case for all models with low He shell masses.
@@
The amount of unburnt He, and its range of ejection velocity in the double detonation scenario can show significant variation for different lines of sight and between models (see e.g. figure 1 in Collins et al. 2022). Explosion simulations for models in which the He detonation is assumed to be ignited through compressional heating of accreted material predict very little unburnt He is present at velocities below . In another widely discussed double-detonation scenario, the initial He shell detonation is instead ignited during the dynamical merger of two CO WDs which leads to a core detonation of the primary WD (Guillochon et al., 2010; Pakmor et al., 2013; Tanikawa et al., 2018, 2019; Boos et al., 2021; Shen et al., 2021b; Pakmor et al., 2022; Boos et al., 2024; Pollin et al., 2024). In this scenario the secondary WD may also explode via a double detonation (Tanikawa et al., 2019; Pakmor et al., 2022; Boos et al., 2024). Explosion simulations of such models predict unburnt He at generally lower velocities compared to accretion ignited double detonation models and for some lines of sight also show significant amounts of unburnt He at velocities below (see e.g. figure 1 of Pollin et al. 2024). The model presented by C23 is representative of a realisation of the double detonation scenario with He present at relatively high velocities (the He distribution peaks at ) and a reasonably large mass of present in the He shell detonation ash. In this paper, we instead investigate the potential of He spectral features to form for a realisation of the double detonation scenario with unburnt He present at significantly lower velocities (the He distribution peaks at ) and only a small mass of present in the He shell detonation ash. In particular, comparing the He spectral features predicted by our new model to those predicted by the model of C23 will provide insight into how sensitive the spectral signatures of He are to different model ejecta structures.
@@
We describe our model ejecta structure and the radiative transfer simulation set up in Section 2. We then present the simulated spectra and their comparisons with SN 2011fe and SN 2022xkq in Section 3 before discussing the implications of the evolution of specific spectral features in Section 4 and presenting our conclusions in Section 5.
@@ 2.2 Double detonation ejecta model
The ejecta model simulated in this study has a similar mass to the C23 ejecta model (0.47 and 0.49 respectively). However, our ejecta model has a total He mass (0.04 ) which is more than double that of the C23 model (0.018 ). Additionally, the ejecta structures show significant differences (compare Figure 1 and figure 1 in C23). In particular, for the C23 model the He lies at a relatively high velocity (peaking at ) and there is a significant mass of ( ) present in the He shell. In contrast, the ejecta model investigated in this study shows a substantial mass of unburnt He at significantly lower velocities (peaking at ) with only a very small mass of (< ) present in this region (such low to He ratios are characteristic of the negative region in the Gronow et al. (2021) models, see Collins et al. 2022 figure 1). As a result our new model has significantly less co-spatial with the unburnt He in the He shell relative to the C23 model. Therefore, the radiative transfer simulation presented here allows us to explore a significantly different realisation of the double detonation scenario than C23.
@@ 3.2.1 Suitability of comparison objects
In the following Sections we focus on the He lines which exhibit the strongest contribution to our model spectra and their comparisons with the spectra of SN 2011fe and SN 2022xkq.
@@ 3.2.2 Simulated He i 10830 Å feature compared to observations
First comparing to SN 2011fe (left panel Figure 2), we see the He feature predicted by our simulation overlaps with the feature Hsiao et al. (2013) identify as Mg ii 10927 Å in SN 2011fe. In the spectra of SN 2011fe this feature shows an extended blue wing which Hsiao et al. (2013) attribute to C i 10693 Å (Marion et al. 2015 propose the same identification for SN 2014J, another normal SN Ia). However, our model does not predict any clear feature from C i 10693 Å (see Section 4.2 for further discussion). The strongest absorption of the model He feature is to the blue wing of the Mg feature in SN 2011fe. We therefore suggest this extended wing may instead be an observational signature of He i 10830 Å. In this interpretation the observed absorption feature forms from a strong blend of Mg ii 10927 Å and He i 10830 Å with He i 10830 Å primarily contributing to the blue wing of the feature. We do however note the He feature predicted by our model (which represents a realisation of the double detonation scenario with a fairly substantial ejected He mass) is significantly too strong to be consistent with SN 2011fe (or SN 2014J). Therefore, further simulations for models with somewhat lower He ejected masses than the model presented here are desirable to investigate if they predict a He contribution to the blue wing of the Mg feature consistent with observations. We also note Mg ii 10927 Å contributes to our model spectrum at wavelengths consistent with the Mg feature observed in the spectrum of SN 2011fe. However the contribution of Mg ii 10927 Å is weaker than that of He i 10830 Å in our simulation for all epochs simulated.
@@
However, SN 2022xkq also shows a weaker secondary absorption feature to the blue of the Mg feature that disappears by peak (this feature is also observed at early times in other transitional SNe Ia, see e.g. Hsiao et al. 2015; Wyatt et al. 2021). Our model does not predict any such double absorption feature and the He i 10830 Å feature predicted by our model forms at too low a velocity to match the secondary absorption feature. However, the double detonation model of C23 does predict such a double feature with Mg ii 10927 Å forming the primary absorption feature and high velocity He i 10830 Å forming the secondary feature. C23 demonstrated this simulated feature was consistent with the double absorption feature observed in the spectrum of another transitional SN, iPTF13ebh (Hsiao et al., 2015) which is very similar to the feature observed in the spectrum of SN 2022xkq. We note that previous studies have suggested this secondary absorption feature observed in the spectra of some SNe Ia is instead formed by C i 10693 Å (e.g. Höflich et al. 2002; Hsiao et al. 2015; Wyatt et al. 2021; Li et al. 2022; Pearson et al. 2024), however neither our model nor the C23 model show any such C i 10693 Å spectral feature (see Section 4.2 for further discussion).
@@ 3.2.3 Optical He spectral contribution and observations
The simulation showed no clearly identifiable He features in the optical, but nonetheless we now discuss how the optical regions with slight He contributions compare with the observations.
@@ 3.2.4 He i 2 m feature and observations
A clear He i 2 m feature is not predicted by this simulation. The 0.04 of He ejected by the model we investigate here is likely towards the upper limit of the ejected He mass that can be present in a model for it to still produce reasonable agreement with the optical spectra of normal and transitional SNe Ia (Kromer et al., 2010). The weak contribution of He i 2 m in our simulation (see Figure 2) therefore suggests it is very unlikely that such a feature would be observable for any normal or transitional SNe Ia. This is consistent with the fact that neither SN 2011fe or SN 2022xkq show any spectral feature around 2 m. We note however that even if this feature can form, observing it will be challenging for ground based telescopes because the spectral contribution of the He i 2 m line predicted by the model lies very close to a telluric region (see right panel in Figure 2) and thus the observed spectra are likely still impacted by the proximity to this region. For example, the observed SN 2022xkq spectrum at 0.1 d after B-peak shows what appears to be an absorption feature around 2 m. However, the potential feature does not show a consistent evolution in the observed spectra and only appears for a single epoch and thus is very unlikely to be a real feature (Pearson et al., 2024). Therefore space based (e.g JWST) observations will likely be required to observe any potential 2 m feature.
4 Discussion
As discussed in Section 2.2 the model presented here represents a double detonation structure with a relatively substantial mass of ejected He (0.04 ), a significant fraction of He present at reasonably low velocities and only a small mass of present in the He shell detonation ash. In contrast the double detonation model presented by C23 has a significantly lower mass of ejected He (0.018 ), He present at higher velocities and a reasonably large mass of present in the He shell detonation ash. In the following Section we first discuss the differences in the He features predicted by each of these models and comment on the implications of these differences for the double detonation scenario. We then discuss the spectral contribution of C in the double detonation models and comment on how this compares with observations.
4.1 Sensitivity of simulated He features to He ejecta structure
As already discussed in Section 3.2.2 the He i 10830 Å feature predicted by the model of C23 shows significant differences compared to the feature predicted by our model. In particular the He feature appears at higher velocities and forms a separate feature to the blue wing of the Mg ii 10927 Å feature. Additionally the He feature in the C23 model evolves to lower velocities with time: at 5 d after explosion the C23 feature is formed at but forms at at 2 weeks after explosion. The inner velocity of the region of unburnt He in the C23 model is at (see C23 figure 1) which suggests the C23 He feature does not depend so strongly on unburnt He close to the inner boundary of the He shell. This is consistent with the distribution of unburnt He in the C23 model which peaks at and shows a much more gradual drop off in density at higher velocities compared to the model presented in this study.
Overall the spectral contribution of He i 10830 Å, and its evolution with phase, strongly depends on the distribution of unburnt He in the model. Given that this can vary significantly for different lines-of-sight and between different double detonation models, we conclude that significant variation in the role of He i in shaping this feature is to be expected.
4.2 Spectral contribution of C in simulations
@@
C i 10693 Å has been proposed to have an observable contribution to the blue wing of the Mg ii 10927 Å feature in SNe Ia before peak, either as an extended blue wing (e.g. Hsiao et al. 2013; Marion et al. 2015) or a distinct secondary feature (e.g. Höflich et al. 2002; Hsiao et al. 2015; Wyatt et al. 2021; Li et al. 2022; Pearson et al. 2024). We therefore comment here on the spectral contribution of C in our simulation.
@@
Our model does show some contribution from C ii in the optical. In particular, a weak C ii 6580 Å feature forms on the red shoulder of the Si ii 6355 Å feature from 7 d before up until B-peak. This feature appears at a velocity generally consistent with the C ii 6580 Å feature observed in the spectra of SN 2011fe and SN 2022xkq. We also see a weak spectral contribution from C ii 7235 Å (again see Figure 3), however the contribution is so weak that it never produces any clear spectral feature.

### v2

@@ Abstract
Models of sub-Chandrasekhar mass double detonations for Type Ia supernovae (SNe Ia) suggest a distinguishing property of this scenario is unburnt helium in the outer ejecta. However, modern explosion simulations suggest there may be significant variations in its mass and velocity distribution. We recently presented a NLTE (non local thermodynamic equilibrium) radiative transfer simulation for one realisation of the double detonation scenario with a modest He mass (0.018 ) present in the ejecta at relatively high velocities (). That simulation predicted a He i 10830 Å feature blueward of Mg ii 10927 Å consistent with near-infrared observations of “transitional” SNe Ia. To demonstrate the expected diversity in the helium signature, here we present a calculation for a double detonation model with a higher He mass (0.04 ) ejected at lower velocities (). Despite our simulation predicting no clear optical or 2 m helium features, a strong and persistent He i 10830 Å absorption is present. The feature appears at wavelengths consistent with the extended blue wing of the Mg ii 10927 Å feature sometimes present in observations, suggesting this is a helium spectral signature (although for this particular model it is too strong and persistent to be consistent with normal SNe Ia). The significant differences in He i 10830 Å predicted by the two simulations suggests helium spectral signatures likely show significant variation throughout the SNe Ia population. This motivates further work to use this observable signature to test the parameter space for double detonation models.
@@
Type Ia supernovae (SNe Ia) are the thermonuclear explosions of at least one white dwarf (WD) in a close binary system. However, the nature of the progenitor systems and explosion mechanisms that produce SNe Ia are still uncertain. The double detonation is a widely discussed sub-Chandrasekhar (sub-) explosion mechanism which has the potential to explain SNe Ia with a variety of luminosities. This scenario involves the ignition of a He detonation in a surface He layer (shell) on a carbon-oxygen (CO) WD (see, e.g., the study of He ignition by Glasner et al., 2018). The shock from the initial He detonation then results in the ignition of a secondary core detonation that entirely disrupts the WD. The mass of He required for a He detonation to occur depends on the specific explosion mechanism. For spontaneous ignition following stable mass transfer from a He star, massive He shells (> ) are required (e.g. Nomoto 1980, 1982; Taam 1980; Livne 1990; Woosley & Weaver 1994; Hoeflich & Khokhlov 1996; Nugent et al. 1997). This results in light curves and spectra incompatible with observations of normal SNe Ia due to the over-abundance of iron-group elements in the outer ejecta produced by the He detonation. If the stable mass transfer is instead from a degenerate companion, significantly less massive He shells are required (< Bildsten et al. 2007; Shen & Bildsten 2009; Shen et al. 2010). We note that while it has been demonstrated that if such a He detonation occurs, a secondary core detonation will likely follow (Fink et al., 2010; Gronow et al., 2020; Gronow et al., 2021; Boos et al., 2021), the initiation of the He detonation may be inhibited when rapid rotation is considered (see e.g. Neunteufel et al. 2017, 2019; Piersanti et al. 2024). In an alternative scenario, the He detonation may be triggered dynamically during the merger of two WDs (Shen & Moore, 2014; Tanikawa et al., 2018, 2019; Pakmor et al., 2022; Shen et al., 2024), requiring only low He shell masses (< 0.1 ). Double detonation simulations invoking low He shell masses predict light curves and spectra that are much more consistent with normal SNe Ia (e.g. Kromer et al. 2010; Woosley & Kasen 2011; Polin et al. 2019; Shen et al. 2021a; Pollin et al. 2024; Collins et al. 2025 although see also Collins et al. 2022).
@@
The amount of unburnt He, and its range of ejection velocity in the double detonation scenario can show significant variation for different lines of sight and between models (see e.g. figure 1 in Collins et al. 2022). Explosion simulations for models in which the He detonation is assumed to be ignited through compressional heating of accreted material predict very little unburnt He is present at velocities below . In another widely discussed double-detonation scenario, the initial He shell detonation is instead ignited during the dynamical merger of two CO WDs which leads to a core detonation of the primary WD (Guillochon et al., 2010; Pakmor et al., 2013; Pakmor et al., 2022; Tanikawa et al., 2018, 2019; Boos et al., 2021; Boos et al., 2024; Shen et al., 2021b; Pollin et al., 2024). In this scenario the secondary WD may also explode via a double detonation (Tanikawa et al., 2019; Pakmor et al., 2022; Boos et al., 2024). Explosion simulations of such models predict unburnt He at generally lower velocities compared to accretion ignited double detonation models and for some lines of sight also show significant amounts of unburnt He at velocities below (see e.g. figure 1 of Pollin et al. 2024). The model presented by C23 is representative of a realisation of the double detonation scenario with He present at relatively high velocities (the He distribution peaks at ) and a reasonably large mass of present in the He shell detonation ash. In this paper, we instead investigate the potential of He spectral features to form for a realisation of the double detonation scenario with unburnt He present at significantly lower velocities (the He distribution peaks at ) and only a small mass of present in the He shell detonation ash. In particular, comparing the He spectral features predicted by our new model to those predicted by the model of C23 will provide insight into how sensitive the spectral signatures of He are to different model ejecta structures.
@@ 2.2 Double detonation ejecta model
The ejecta model simulated in this study is based on the negative (south pole) region of the 3D M08_03 explosion model Gronow et al. (2020) presented (the He detonation is ignited on the positive axis) while the C23 model is based on an equatorial slice of the 3D M2a explosion model presented by Gronow et al. (2020). As such the models show significant differences in their ejecta: although our model has a mass similar to the C23 ejecta model (0.47 and 0.49 respectively) it has a total He mass (0.04 ) which is more than double that of the C23 model (0.018 ). Additionally, the ejecta structures show significant differences (compare Figure 1 and figure 1 in C23). In particular, for the C23 model the He lies at a relatively high velocity (peaking at ) and there is a significant mass of ( ) present in the He shell. In contrast, the ejecta model investigated in this study shows a substantial mass of unburnt He at significantly lower velocities (peaking at ) with only a very small mass of (< ) present in this region (such low to He ratios are characteristic of the negative region in the Gronow et al. (2021) models, see Collins et al. 2022 figure 1). As a result our new model has significantly less co-spatial with the unburnt He in the He shell relative to the C23 model. Therefore, the radiative transfer simulation presented here allows us to explore a significantly different realisation of the double detonation scenario than C23.
@@ 3.1.3 NIR He i 2 m
3.2 Comparisons to C23 and implications for He signature diversity
As discussed in Section 2.2 the model presented here represents a double detonation structure with a relatively substantial mass of ejected He (0.04 ), a significant fraction of He present at reasonably low velocities and only a small mass of present in the He shell detonation ash. In contrast the double detonation model presented by C23 has a significantly lower mass of ejected He (0.018 ), He present at higher velocities and a reasonably large mass of present in the He shell detonation ash. In the following we discuss the differences in the He features predicted by each of these models and comment on the implications of these differences for the double detonation scenario.
In contrast, the He i 10830 Å feature predicted by the model of C23 appears at higher velocities and forms a separate feature to the blue wing of the Mg ii 10927 Å feature (see C23 figure 3). The He feature in the C23 model also evolves to lower velocities with time: at 5 d after explosion the C23 feature is formed at but forms at at 2 weeks after explosion. The inner velocity of the region of unburnt He in the C23 model is at (see C23 figure 1) which suggests the C23 He feature does not depend so strongly on unburnt He close to the inner boundary of the He shell. This is consistent with the distribution of unburnt He in the C23 model which peaks at and shows a much more gradual drop off in density at higher velocities compared to the model presented in this study.
Neither our model nor the model of C23 predict any clear optical He i features. Additionally, while the C23 simulation did not have sufficient signal to noise to determine if a m feature appears, our simulation predicts no such feature forms, despite the much stronger and more persistent He i 10830 Å absorption relative to the C23 simulation. Together these models therefore suggest that He i 10830 Å is the most robust spectral signature of He and for double detonation models with moderate He shell masses (such as our model and the C23 model) other He i spectral lines (either optical or m) that can act as corroborating signatures will not necessarily appear. As such the absence of these He features should not be used to rule out the identification of He i 10830 Å.
Comparing the predictions of the two models also demonstrates that the spectral contribution of He i 10830 Å and its evolution with phase, strongly depends on the distribution of unburnt He in the model. Given that this can show substantial differences between models and with line-of-sight, we conclude that significant variation in the role of He i in shaping this feature is to be expected. This suggests there is substantial diversity in the observed He i 10830 Å spectral signature throughout the SN Ia population.
@@ 3.3.2 Simulated He i 10830 Å feature compared to observations
First comparing to SN 2011fe (left panel Figure 2), we see the He feature predicted by our simulation overlaps with the feature Hsiao et al. (2013) identify as Mg ii 10927 Å in SN 2011fe. In the spectra of SN 2011fe this feature shows an extended blue wing which Hsiao et al. (2013) attribute to C i 10693 Å (Marion et al. 2015 propose the same identification for SN 2014J, another normal SN Ia). However, our model does not predict any clear feature from C i 10693 Å (see Section 3.3.5 for further discussion). The strongest absorption of the model He feature is to the blue wing of the Mg feature in SN 2011fe. We therefore suggest this extended wing may instead be an observational signature of He i 10830 Å. In this interpretation the observed absorption feature forms from a strong blend of Mg ii 10927 Å and He i 10830 Å with He i 10830 Å primarily contributing to the blue wing of the feature. We do however note the He feature predicted by our model (which represents a realisation of the double detonation scenario with a fairly substantial ejected He mass) is significantly too strong to be consistent with SN 2011fe (or SN 2014J). Therefore, further simulations for models with somewhat lower ejected He masses than the model presented here are desirable to investigate if they predict a He contribution to the blue wing of the Mg feature consistent with observations. We also note Mg ii 10927 Å contributes to our model spectrum at wavelengths consistent with the Mg feature observed in the spectrum of SN 2011fe. However the contribution of Mg ii 10927 Å is weaker than that of He i 10830 Å in our simulation for all epochs simulated.
@@
However, SN 2022xkq also shows a weaker secondary absorption feature to the blue of the Mg feature that disappears by peak (this feature is also observed at early times in other transitional SNe Ia, see e.g. Hsiao et al. 2015; Wyatt et al. 2021). Our model does not predict any such double absorption feature and the He i 10830 Å feature predicted by our model forms at too low a velocity to match the secondary absorption feature. However, as discussed in Section 3.2 the double detonation model of C23 does predict such a double feature with Mg ii 10927 Å forming the primary absorption feature and high velocity He i 10830 Å forming the secondary feature. C23 demonstrated this simulated feature was consistent with the double absorption feature observed in the spectrum of another transitional SN, iPTF13ebh (Hsiao et al., 2015) which is very similar to the feature observed in the spectrum of SN 2022xkq. We note that previous studies have suggested this secondary absorption feature observed in the spectra of some SNe Ia is instead formed by C i 10693 Å (e.g. Höflich et al. 2002; Hsiao et al. 2015; Wyatt et al. 2021; Li et al. 2022; Pearson et al. 2024), however neither our model nor the C23 model show any such C i 10693 Å spectral feature (see Section 3.3.5 for further discussion).
@@ 3.3.3 Optical He spectral contribution and observations
The simulation showed no clearly identifiable He features in the optical (as was the case for the model of C23), but nonetheless we now discuss how the optical regions with slight He contributions compare with the observations.
@@ 3.3.4 He i 2 m feature and observations
A clear He i 2 m feature is not predicted by this simulation. The 0.04 of He ejected by the model we investigate here is likely towards the upper limit of the ejected He mass that can be present in a model for it to still produce reasonable agreement with the optical spectra of normal and transitional SNe Ia (Kromer et al., 2010). The weak contribution of He i 2 m in our simulation (see Figure 2) therefore suggests it is very unlikely that such a feature would be observable for any normal or transitional SNe Ia. This is consistent with the fact that neither SN 2011fe or SN 2022xkq show any spectral feature around 2 m. Therefore, the presence of strong He i 10830 Å does not require that a 2 m feature is also observed. We note however that even if this feature can form, observing it will be challenging for ground based telescopes because the spectral contribution of the He i 2 m line predicted by the model lies very close to a telluric region (see right panel in Figure 2) and thus the observed spectra are likely still impacted by the proximity to this region. For example, the observed SN 2022xkq spectrum at 0.1 d after B-peak shows what appears to be an absorption feature around 2 m. However, the potential feature does not show a consistent evolution in the observed spectra and only appears for a single epoch and thus is very unlikely to be a real feature (Pearson et al., 2024). Therefore space based (e.g JWST) observations will likely be required to observe any potential 2 m feature.
3.3.5 Spectral contribution of C and observations
@@
C i 10693 Å has been proposed to have an observable contribution to the blue wing of the Mg ii 10927 Å feature in SNe Ia before peak, either as an extended blue wing (e.g. Hsiao et al. 2013; Marion et al. 2015) or a distinct secondary feature (e.g. Höflich et al. 2002; Hsiao et al. 2015; Wyatt et al. 2021; Li et al. 2022; Pearson et al. 2024). We therefore comment here on the spectral contribution of C in our simulation and the C23 simulation.
@@
Our model does show some contribution from C ii in the optical. In particular, a weak C ii 6580 Å feature forms on the red shoulder of the Si ii 6355 Å feature from 7 d before up until B-peak. This feature appears at a velocity generally consistent with the C ii 6580 Å feature observed in the spectra of SN 2011fe and SN 2022xkq. We also see a weak spectral contribution from C ii 7235 Å (again see Figure 3), however the contribution is so weak that it never produces any clear spectral feature. We note that while the C23 simulation also shows contributions from C ii 6580 and 7235 Å, unlike our simulation no clear C ii 6580 Å feature forms due to blending with stronger lines.